How many grams of 

KClO3 must be decomposed to produce 91.7 g O


2

91.7 g O2?


1
Expert's answer
2022-01-30T07:58:34-0500

Solution:

Calculate the moles of O2:

The molar mass of O2 is 32.0 g/mol

Hence,

Moles of O2 = (91.7 g O2) × (1 mol O2 / 32.0 g O2) = 2.8656 mol O2


Balanced chemical equation:

2KClO3 → 2KCl + 3O2

According to stoichiometry:

2 mol of KClO3 produces 3 mol of O2

X mol of KClO3 produces 2.8656 mol O2:

Thus,

Moles of KClO3 = X = ( 2.8656 mol O2) × (2 mol KClO3 / 3 mol O2) = 1.91 mol KClO3


Calculate the mass of KClO3:

The molar mass of KClO3 is 122.55 g/mol

Hence,

Mass of KClO3 = (1.91 mol KClO3) × (122.55 g KClO3 / 1 mol KClO3) = 234.07 g KClO3 = 234.1 g KClO3

Mass of KClO3 is 234.1 grams


Answer: 234.1 grams of KClO3 must be decomposed.